Use the following balanced equation to answer the parts below:


3H2​ + N2 ​ 2NH3

How many grams of NH3 can be made from 5.60  mol of H2 and excess N2?

How many grams of NH3 can be made from 4.34 mol of N2 and excess H2?

If you want to make 19.59 mol of NH3, how many grams of N2 do you need?




1
Expert's answer
2021-04-11T23:50:36-0400

Molar Mass of NH3 = 17.03

Molar Mass of N2 = 28.0134

Molar Mass of H2 = 2.016

How many grams of NH3 can be made from 5.60  mol of H2 and excess N2?

From the equation

= 3/4×5.6 = 4.2moles of H2

= 1/4 × 5.6 = 1.4 moles of N2

3moles of H2 = 2 moles of NH3

2/3 × 4.2 = 2.8 moles of NH3

= 2.8 × 17.031 = 47.6868g

1 mole of N2 = 2 moles of NH3

= 1.4 × 2 = 2.8 × 17.031= 47.6868g

= 47.6868 + 47.6868 = 95.3736g of NH3

How many grams of NH3 can be made from 4.34 mol of N2 and excess H2?

= 3/4×4.34 = 3.255 moles

1/4×4.34 = 1.085moles

= 2/3×3.255 = 2.17 moles of NH3

= 2.17×17.031 = 36.95727g

1.085 × 2 = 2.17

= 36.95727g

= 36.95727×2 = 73.9145g of NH3

If you want to make 19.59 mol of NH3, how many grams of N2 do you need?

19.59 moles of NH3

2 moles of NH3 = 1mole of N2

= 19.59/2 = 9.795× 28.0134

= 274.391253g of N2